10 Points 10, 7 years ago, the City of Calabasas issued $5,000,000 of 10.5% coupon, 20-year annual payment, tax-exempt muni bonds. The bonds had 5 years of call protection, but now the bonds can be called if the city chooses to do so. The call premium would be 6.5% of the face amount. New 13-year, 8.5%, annual payment bonds can be sold at par, but flotation costs on this issue would be 4%, of the amount of bonds sold, what is the net present value of the refunding? Note that cities pay no in are not relevant come taxes, hence taxes
Answer #1
PW of floation costs on new issue =5000000*4% =$ -2,00,000
PW of Call premium on old bonds = 5000000*6.5% =$ -3,25,000
Interest saved per annum = 5000000*(10.5-8.5)% = 100000
PW of interest saved for 13 years at 8.5% = 100000*(1.085^13-1)/(0.085*1.085^13) =$   7,69,095
NPV of refunding$   2,44,095
As the NPV of refunding is positive, the new bonds can be issued and the old bonds can be called.


answered by: ANURANJAN SARSAM